Jerri Kingston bought 100 shares o stock at S86 84 per share, using an nit a margin o 53%. Glven a ma tenance margin o 32%, how ar does he stock have o drop be re Jer aces a margin cal ? Assume that there are no other securit es n the margin account. Before Jerri faces a margin call, the stack has to drop to Sper share (Raund to the nearest cent
Answer #1

Loan Amount = (1-53%)*100*86.84 = 4081.48

Maintenance Margin = (100P - 4081.48)/100P = 0.32

100P - 4081.48 = 32P
78 P = 4081.48
P = 4081.48/78 = 52.33

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well